OpenStudy (anonymous):

A 2.0 μC test charge travels along an equipotential line a distance of 0.50 cm between two parallel charged plates with a field strength of 500.0 N/C. What is the change in voltage? (Points : 5) -100 V +100 V +1,000 V 0 V

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

@IrishBoy123

OpenStudy (michele_laino):

by definition, an equipotential line, is a line whos potential is constant,for example we can consider an equipotential line whose potential is equal to 100 volts. So the requested potential difference is: \[\Delta V = \frac{{\Delta U}}{q} = \frac{{100 - 100}}{q} = ...\]

OpenStudy (anonymous):

0 V